# Insights into LLM Long-Context Failures: When Transformers Know but Don’t Tell

### v1

@@ Appendix B Probing Setup
In the experiment described in §3, we employ linear classifiers as our probing method.
For any given task, we choose $\{1,0.1n,0.2n,\dots,1.0n\}$-th position as gold ID. Following the prompt format in §A, we generate prompts with all chosen IDs, for $10,000$ iterations, resulting in a set of 110,000 prompts.
@@
Each prompt is fed into language model, and the embedding from each layer’s last token is collected. For each layer, separately, we have $110,000$ embeddings corresponding to 11 IDs and train a classifier for ten times, with embedding as input and ID as output. We calculate their mean accuracy and standard deviation.
Appendix C Experiments Results on Mistral-7B-Instruct-v0.3 Jiang et al. (2023) and Gemma-7b-it Team et al. (2024)
@@
We conduct same experiment procedure on additional two models, which produce the same pattern. The experiment is running on one A100 GPU.
Figure 5: In the left figure, Mistral’s probing result in 100 kv-pairs retrieval task resembles with Llama3’s in most ways, but there’re still a few differences. First, the trends of locating information from the head and end are more similar compared to those in Llama3. Also the model takes more layers than Llama3 to well locate information in middle context. The right figure highlights the significant discrepancy between the probing peak accuracy and generation accuracy, indicating a severe ’know don’t tell’ phenomenon.
Figure 6: Mistral’s layer-wise probing classifier accuracy performs in the same pattern as Llama3. Middle context takes more layers to encode and results in a lower peak accuracy. In right figure, both its peak accuracy and generation accuracy shows a U-shape curve, with still probing consistently outperforms generation.
Figure 7: In the left figure, Gemma shows a similar pattern like other models in 100 kv-pairs retrieval task. However, information from all positions is located more slowly than other models: Information in the head or end position of input context is located after 5 layers, while information in the middle of input takes 15 layers to be located. In the right figure, it shows a significant gap between generation accuracy and probing peak accuracy, which is similar with Mistral. This highlights a significant ’know don’t tell’ phenomenon.
Figure 8: For Gemma MDQA task, although head and end context follows same observation, it middle context shows a sudden decrease in accuracy, indicating a sudden information lost but soon retrieve it back. The right figure follows the same pattern, where generation accuracy is consistently low, disconnecting from the high U-shape probing accuracy curve.

### v2

@@ Appendix B Probing Setup
In the experiment described in §3, we employ linear classifiers as our probing method. For any given task, we choose $\{1,0.1n,0.2n,\dots,1.0n\}$-th position as gold ID, where $n$ is the number of documents ($0.1n$ for $n=30$ means at ID $3$). Following the prompt format in §A, we generate prompts with all chosen IDs, for $10,000$ iterations, resulting in a set of 110,000 prompts. Each prompt is fed into language model, and the embedding from each layer’s last token is collected. For each layer, separately, we have $110,000$ embeddings corresponding to 11 IDs and train a classifier for ten times, with embedding as input and ID as output. We calculate their mean accuracy and standard deviation.
Appendix C Experiments Results on Gemma-7b-it Team et al. (2024)
@@
Following §4.1 and §4.2, we conduct same experiment procedure on one additional model, which produces the same pattern. The experiment is running on one A100 GPU.
Figure 5: Replicating the results of Fig.2 and Fig.3 using the Gemma model with 100 kv-pairs. The findings for this model also align with the observations in §4.1 and §4.2. On the right, there is a notable gap between generation accuracy and peak probing accuracy, mirroring the results observed with Mistral in the main text.
Figure 6: Replicating the results of Fig.2 and Fig.3 using the Gemma model with 30-document MDQA. The findings for this model also align with the observations in §4.1 and §4.2. On the left, while the beginning and end contexts follow the same pattern, the middle context exhibits a sudden drop in accuracy, indicating a brief loss of information, which is quickly regained.
Appendix D 2D Visualization of Hidden States Per Layer
We illustrate the model’s internal representations by employing Principal Component Analysis (PCA) (Shlens, 2014) (Fig. 7) and Uniform Manifold Approximation and Projection (UMAP) (McInnes et al., 2020) (Fig. 8) to reduce the hidden states to two dimensions. Each subplot depicts the hidden states of a single layer within the Transformer. The subplots feature 11 colored points, each representing a different gold document ID. The color-to-ID mapping can be found in the color bar located in the upper right corner. We reuse the 30-documents prompts. Among this same set of documents, we rotate the ID of the gold document to 11 equally separated positions, as in §B. For the 11 prompts, we extract the last token dimensions from LLaMa3-8B-Instruct and perform dimensional reduction. Each dots in the plot represent a single data point.
Figure 7: PCA visualization across layers: We apply PCA dimension reduction to the last embedding of each layer and visualize the results. The color of the dots indicates the position of the gold document in the prompt, ranging from 0% (gold document at the beginning) to 100% (gold document at the end). We observe that the dots are entangled in the early layers, start to form a path in the order of gold document ID in the middle layers, and this order diminishes in the later layers. This observation further supports our findings in §4.2, indicating that the internal representation of LLMs becomes richer with more layers, although the final layer may not exhibit the peak of representation richness.
Figure 8: UMAP visualization across layers reveals a pattern similar to that in PCA Fig. 7. The embeddings become less entangled in the middle layers and appear more random in the extreme (earlier and final) layers.
Below, we calculate the average distance between points in the PCA visualization (Fig.7). For each layer $l$ (each subfigure), there are 11 points ${x^{l}_{1},x^{l}_{2},x^{l}_{3},\dots,x^{l}_{11}}$, each corresponding to a different gold ID. We compute the average distance for layer $l$ as $d_{l}=\sum_{i=1}^{10}\text{distance}(x^{l}_{i},x^{l}_{i+1})/10$. As shown in Fig.9, the average distance $d_{l}$ increases up to layer 21, indicating a growing degree of separability. After reaching the peak distance, it decreases until the final layer. This pattern further supports our probing results. Additionally, as observed in §4.3, most gold IDs achieve peak accuracy around layer 18, which is very close to the peak distance layer (layer 20) identified here.
Figure 9: Average distance between scatter points across layers. The average distance has an obvious increasing trend during early layers, after achieves peak value around layer 21, it gradually decreases. The overall trend is similar to the probing accuracy in Fig. 3.
Appendix E Language Generation Logits Trends Across Layers
We apply the logit lens (nostalgebraist, 2021) to examine the layers where the model begins to produce more probability mass over the correct answer. Specifically, for each layer, we multiply the last token embedding $x^{l}$ with the LM Head and apply SoftMax to obtain the generation distribution over the entire vocabulary. We then record the probability of generating the correct token. The logit lens enables internal analysis of the LLM output distribution flow across layers. We reuse the 30-document prompts and apply the logit lens to each. Fig. 10 shows how the logits for the correct answer evolve across layers, with visualizations for prompts containing the gold document at different locations.
Figure 10: Logits of generating the first token of the correct answer across layers. The lines indicate different gold document ID. x-axis is the layer index, and y-axis is the logits across 1000 different prompts. During middle layers, blue line(ID 30) rises much more significantly than other IDs, however it also drops as the lowest at the last layer. For other lines, they share similar patterns in most layers but diverge in last few layers.
Layer-wise logit divergence.
The changes in logits across layers reveal certain key observations:
  1. 1.
    In all prompts, the logits follow an almost identical pattern before layer 20, showing a steady, minimal increase. The divergence begins at layer 20.
  2. 2.
    At layer 20, when the gold document is placed at the end of the prompt, there is a sharp and significant increase in the logits. This is supported by Fig. 3, where the same prompt achieves near-perfect probing accuracy.
  3. 3.
    All prompts reach their peak logit at layer 31, but exhibit a positional bias. The later the gold document appears in the prompt, the higher the peak logit.
These findings reveal that positional bias emerges during the model’s internal processing, with a noticeable shift at a certain layer, where the position of gold information influences the quality of language output.